### v1

# Self-supervised Physics-guided Model with Implicit Representation Regularization for Fast MRI Reconstruction

Abstract
Magnetic Resonance Imaging (MRI) is a vital clinical diagnostic tool, yet its widespread application is limited by prolonged scan times. Fast MRI reconstruction techniques effectively reduce acquisition duration by reconstructing high-fidelity MR images from undersampled k-space data. In recent years, deep learning-based methods have demonstrated remarkable progress in this field, with self-supervised and unsupervised learning approaches proving particularly valuable in scenarios where fully sampled data are difficult to obtain. This paper proposes a novel zero-shot self-supervised reconstruction framework named UnrollINR, which enables scan-specific MRI reconstruction without relying on external training data. The method adopts a physics-guided unrolled iterative reconstruction architecture and introduces Implicit Neural Representation (INR) as a regularization prior to effectively constrain the solution space. By combining a deep unrolled structure with the powerful implicit representation capability of INR, the model’s interpretability and reconstruction performance are enhanced. Experimental results demonstrate that even at a high acceleration rate of 10, UnrollINR achieves superior reconstruction performance compared to the supervised learning method, validating the superiority of the proposed method.
{IEEEkeywords}
Implicit neural representation, fast MRI, parallel imaging, scan-specific, self-supervised learning
1 Introduction
[…]
Magnetic Resonance Imaging (MRI) is a safe, non-invasive, radiation-free imaging technique with excellent soft tissue contrast, making it an indispensable tool in clinical diagnostics. However, prolonged scanning times remain a major technical limitation [71]. In practical applications, undersampling below the Nyquist rate is commonly employed to accelerate MRI by reducing $k$-space data acquisition. Nevertheless, directly applying the inverse Fourier transform to undersampled data leads to spectral aliasing artifacts.
Fast MRI reconstruction is fundamentally a class of regularized inverse problems aimed at recovering high-quality, artifact-free images from undersampled data [104]. Currently, one of the most widely used techniques is parallel imaging (PI), which leverages redundancy in spatial sensitivity from multi-channel receiver coils to effectively suppress artifacts through joint reconstruction [67, 89, 80]. Another commonly employed technique is compressed sensing (CS), which exploits the compressible nature of MRI by employing pseudo-random sparse sampling combined with sparsity or low-rank constraints to achieve accurate reconstruction from undersampled data [79]. However, both PI and CS still face limitations in achieving higher acceleration rates [92, 93].
In recent years, deep learning techniques have demonstrated substantial progress in fast MRI reconstruction. Existing approaches can be broadly categorized into two classes: purely data-driven methods and physics-guided methods. Purely data-driven methods aim to learn an end-to-end mapping from undersampled k-space (or artifact-corrupted images) to fully-sampled $k$-space (or artifact-free images) directly [74], with representative examples including AUTOMAP [110] and U-Net [72, 70]. Although these methods have shown promising reconstruction performance, their heavy reliance on data-driven learning often results in limited model interpretability. In contrast, physics-guided methods seek to integrate model-based image fidelity constraints with deep learning-based data-driven regularization, explicitly incorporating the physical principles of the MRI data acquisition process into the neural network architecture [69]. These approaches are typically implemented by unrolling iterative reconstruction algorithms into networks with a fixed number of steps, where each step alternates between data consistency updates and neural network-based regularization [84]. Representative methods in this category include ADMM-Net [98], ISTA-Net [108], MoDL [58], and VarNet [97]. Such deep unrolled models not only preserve algorithmic interpretability but also significantly enhance reconstruction performance, thereby effectively bridging conventional iterative reconstruction techniques with modern deep learning frameworks.
Typically, deep learning-based MRI reconstruction methods rely on supervised learning strategies, which require large-scale paired training data consisting of fully-sampled and undersampled measurements [85]. However, this approach faces two major challenges in practical applications: first, the high cost of data acquisition makes it difficult and expensive to build large-scale, high-quality paired datasets; second, the issue of out-of-domain generalization arises, where model performance degrades significantly when the test data distribution differs from that of the training set [60]. Consequently, unsupervised learning has emerged as a promising paradigm that does not depend on external paired data [90, 59]. Instead, it aims to mitigate data scarcity by learning the inherent distribution characteristics of the undersampled data itself. As an important branch of unsupervised learning, self-supervised learning further leverages the intrinsic structures or constraints within MRI data as supervisory signals to guide model optimization, thereby improving both reconstruction quality and data utilization efficiency [106, 105].
Implicit Neural Representation (INR), as an emerging self-supervised approach, has recently become an effective tool for solving imaging inverse problems [83, 112]. This method does not rely on external data; instead, it utilizes a neural network, typically a multi-layer perceptron (MLP), to model image intensity as a continuous function of spatial coordinates, thereby achieving a flexible and continuous representation of images [65]. Thanks to this continuous representation capability, INR can more effectively capture fine details and complex structures in images compared to traditional discrete grid-based neural networks. Moreover, the inherent bias of INR can be regarded as an effective form of implicit regularization, which helps stabilize the solution of inverse problems [91].
Currently, INR has demonstrated considerable potential in fast MRI reconstruction. For instance, NeRP is an INR-based method that leverages prior embedding, enabling MRI reconstruction from radially undersampled k-space data [94]. However, since it requires a fully-sampled prior image from the same subject, this approach is primarily suitable for long-term longitudinal imaging scenarios. IMJENSE integrates INR with PI to achieve joint estimation of coil sensitivity maps and images [66]. Both NeRP and IMJENSE are designed for two-dimensional fast MRI reconstruction tasks. Furthermore, INR-based approaches have been widely applied in dynamic MRI reconstruction and parametric quantitative MRI. In the domain of dynamic imaging, representative methods include CineJENSE [68], FMLP [75], and IMJ-PLUS [95], while in parametric imaging, algorithms such as INR-QSM [109] and SUMMIT [76] have been developed. Nevertheless, most existing INR methods predominantly rely on the implicit regularization capability of the network itself. Under high acceleration rates, these methods often exhibit instability, which limits their reconstruction performance.
This study proposes UnrollINR, a novel zero-shot self-supervised reconstruction framework. This approach enables high-quality, scan-specific image reconstruction using only undersampled $k$-space data from a single subject, without requiring any fully-sampled reference data. In physics-guided unrolled iterative reconstruction methods, the design of the neural network-based regularization term plays a crucial role in determining reconstruction performance. The continuous implicit representation capability of INR serves as an effective form of implicit regularization, offering a new approach for solving imaging inverse problems. The core innovation of the proposed unrollINR lies in the deep integration of physics-guided unrolled iterative reconstruction methods with the implicit representation capabilities of INR: the forward physical model ensures data consistency, while the image reconstructed by the INR serves as an effective image prior, embedded within the iterative optimization process. This joint design aims to significantly enhance reconstruction quality and robustness under high acceleration rates, without reliance on external datasets. We evaluated the proposed method on two public datasets and one prospectively undersampled dataset. Experimental results demonstrate that even when trained exclusively on undersampled data, the proposed UnrollINR outperforms the supervised learning approach, confirming its superiority. The main contributions of this work are summarized as follows:
  1. 1.
    We propose UnrollINR, a novel zero-shot self-supervised reconstruction framework for fast MRI reconstruction, which achieves superior performance using only undersampled $k$-space data from a single subject, without relying on external training data.
  2. 2.
    We demonstrate that the reconstruction generated by INR can serve as an effective image prior and an explicit regularization term to enhance the reconstruction model.
  3. 3.
    We integrate a physics-guided unrolled iterative reconstruction algorithm with the INR-based approach, thereby enhancing reconstruction performance while maintaining favorable model interpretability.
  4. 4.
    Extensive validation on publicly available and prospectively undersampled datasets demonstrates that UnrollINR outperforms the supervised learning method.
2 Related Works
2.1 Accelerated MRI Acquisition Model
In multi-channel MRI, multiple receiver coils are utilized to sample frequency-domain raw data, also known as $k$-space. To accelerate the acquisition process, $k$-space measurements are often downsampled via an undersampling mask. Accordingly, the forward model of the multi-coil MRI acquisition can be formulated as follows:
$y_{i}=MFC_{i}x+n_{i},$ (1)
where $x$ denotes the underlying image to be reconstructed, $y_{i}$ is the measured data from the $i$th coil, $n_{i}$ is the corresponding noise, $M$ is the diagonal undersampling mask, $F$ represents the Fourier transform matrix, and $C_{i}$ is the sensitivity map matrix of the $i$th coil. Consequently, the multi-coil acquisition model can be compactly formulated across the coil dimension as
$y=Ex+n,$ (2)
where $E$ represents the forward encoding operator, constructed by concatenating $MFC_{i}$ across the coil dimension. The recovery of $x$ can be mathematically cast as the following optimization problem:
$\arg\min_{x}\left\|y-Ex\right\|_{2}^{2}+\lambda\mathcal{R}({x}),$ (3)
where $\left\|y-Ex\right\|_{2}^{2}$ denotes the data fidelity term that enforces data consistency (DC) with the acquired measurements $y$, $\mathcal{R}({x})$ is the regularization term that enforces prior knowledge and constraints the solution space, and $\lambda$ is a regularization parameter balancing these two terms.
2.2 Deep Unrolled Method
In recent years, deep learning has been widely applied to fast MRI reconstruction. Current deep learning-based MRI reconstruction methods can be broadly categorized into two groups: the first is data-driven end-to-end learning, which aims to directly learn the mapping between undersampled and fully-sampled data; the other is physics-guided unrolled iterative deep learning, which solves the inverse problem of MRI reconstruction by unrolling traditional iterative algorithms into network structures [106]. Compared with data-driven approaches, physics-guided unrolled methods integrate physical priors into the network architecture, offering stronger theoretical guarantees and improved interpretability. These methods incorporate vital prior knowledge through regularization terms during the reconstruction process while maintaining fidelity to the original optimization problem, thereby facilitating a deeper understanding of model behavior and reconstruction results.
Currently, a variety of unrolled iterative methods have been proposed to solve the optimization problem in (3) by leveraging different optimization algorithms, including the alternating direction method of multipliers (ADMM) [63], the iterative shrinkage-thresholding algorithm (ISTA) [61], the proximal gradient descent method [81], and the variable splitting with quadratic penalty [57], among others. In this work, variable splitting with quadratic penalty is employed to reformulate (3) as
$\arg\min_{x,z}\left\|y-Ex\right\|_{2}^{2}+\lambda\left\|x-z\right\|_{2}^{2}+\mathcal{R}({z}),$ (4)
where $z$ is an auxiliary variable which is initially constrained to equal $x$. Hence, the optimization problem in (3) can thus be decomposed into the following two sub-problems as
$z^{t}=\arg\min_{z}\lambda\left\|x^{t-1}-z\right\|_{2}^{2}+\mathcal{R}({z}),$ (5)
$x^{t}=\arg\min_{x}\left\|y-Ex\right\|_{2}^{2}+\lambda\left\|x-z^{t}\right\|_{2}^{2},$ (6)
where $x^{(0)}$ is the initial zero-filled image derived from the undersampled measurements, with $x^{(t)}$ and $z^{(t)}$ being the estimated image and an intermediate variable at the $t$th iteration, respectively. These two problems are addressed alternatively within each iteration. Specifically, the regularization sub-problem in (5) is implictly solved by various neural networks, whereas the DC sub-problem in (6) can be solved using the normal equations
$x^{t}=(E^{H}E+\lambda I)^{-1}(E^{H}y+\lambda z^{t}),$ (7)
where $I$ denotes the identity operator and $(\cdot)^{H}$ denotes the conjugate transpose operator. In practical MRI, the acquired data are often multi-channel, and since $(E^{H}E+\lambda I)^{-1}$ in (7) is not analytically invertible, (7) can be solved using iterative numerical methods such as conjugate gradient (CG) or gradient descent [58]. Existing unrolled reconstruction methods differ primarily in their chosen optimization algorithm for the data consistency sub-problem and the design of the neural network that implicitly solves the regularization sub-problem.
2.3 Implicit Neural Representation
In recent years, Implicit Neural Representations (INRs) have garnered significant attention in the field of computer vision. The core idea is to use a neural network, typically a Multilayer Perceptron (MLP), to model an image or volumetric data as a continuous function of spatial coordinates [96, 88]. Specifically, the network takes spatial coordinates as input and outputs the corresponding image intensity values at those coordinates. Once the network training converges, the continuous representation of the image is implicitly stored within its weights.
The pioneering work on Neural Radiance Fields (NeRF) first successfully applied INRs to 3D scene modeling [82]. By representing a scene as a continuous neural radiance field, it enables the synthesis of novel views with high visual quality. In recent years, the application of INRs in medical imaging has become increasingly widespread, exemplified by techniques such as arbitrary-scale MRI super-resolution [103, 102] and the development of scan-specific reconstruction methods based on INRs, offering new solutions for fast MRI reconstruction, etc. When an INR is employed for fast MRI reconstruction, the underlying network can be formulated as
$f_{\theta}:v=(v_{x},v_{y})\in\mathbb{R}^{2}\rightarrow I\in\mathbb{C},$ (8)
where $v$ represents the coordinate and $I$ denotes the corresponding image intensity.
Meanwhile, related studies have shown that INR networks exhibit significant limitations in representing high-frequency details when processing raw input coordinates directly [91]. To address this issue, various coordinate encoding functions have been proposed. The core idea is to map the low-dimensional input coordinates into a high-dimensional space via a mapping mechanism, thereby enhancing the model’s capacity to capture high-frequency information. These encoding functions can be primarily categorized into two types: fixed encodings and learnable encodings [111]. Fixed encodings, such as positional encoding [82], Fourier feature encoding [100], and radial encoding [78], employ predefined transformation rules. In contrast, learnable encodings introduce learnable parameters and efficient sparse data structures, achieving superior convergence performance and reconstruction accuracy [99, 77, 73]. Based on this, our study adopts the multi-resolution hash encoding technique, a method under the learnable encoding paradigm, to process the input coordinates [87]. And (8) can be rewritten as
$f_{\theta}:\phi(v)\rightarrow I\in\mathbb{C},$ (9)
where $\phi(\cdot)$ represents the coordinate encoding function.
3 Methodology
3.1 Overall Framework
This paper presents a scan-specific self-supervised method named UnrollINR, designed for fast 2D MRI reconstruction. The overall framework is illustrated in Figure. 1. The method adopts a physics-guided unrolled iterative architecture. The reconstruction network consists of several cascading basic units, each comprising a regularization module and a data consistency (DC) module. The regularization module is implemented via an INR-based neural network to exploit prior information within the data. In contrast, the DC module is solved using the conjugate gradient method, ensuring that the reconstruction adheres to the physical acquisition model. The network input, $x^{0}$, is an initial zero-filled image, obtained by applying zero-filling to the original undersampled k-space measurements $y$ followed by an inverse Fourier transform to the image domain, $x^{0}=E^{H}y$. The network output $\widehat{x}$ is the final high-fidelity reconstructed image.
Figure 1: Overview of the proposed UnrollINR framework with an unrolled iterative architecture incorporating an INR-based network as the regularization term.
3.2 Basic Unit of the Proposed Deep Unrolling Network
The inherent implicit learning consistency bias of INRs serves as an effective regularization constraint for solving inverse problems. Therefore, the proposed method, UnrollINR, incorporates an INR-based neural network as the regularization term to exploit image priors within the MRI reconstruction framework. This INR-based regularization term establishes a mapping between MRI spatial coordinates and their corresponding image intensities via a learnable continuous mapping function. As shown in Figure.1, the continuous mapping function $f_{\theta}$ is approximated by an MLP. To enhance the model’s capacity for fitting high-frequency details, the input coordinates $v$ are first mapped into a higher-dimensional feature space using a learnable coordinate encoding function $\phi(\cdot)$ before being fed into the MLP.
This study employs multi-resolution hash encoding as the coordinate encoding function $\phi(\cdot)$ [87]. Compared to other encoding methods, it provides a more compact signal representation. Furthermore, the combination of hash encoding with an MLP enables the use of a shallower MLP to achieve faster convergence and superior detail reconstruction performance. Specifically, the method organizes trainable parameters into $L$ distinct resolution levels, each corresponding to an independent hash table that stores learnable feature vectors. Each level contains $T$ feature vectors of dimension $F$, resulting in $T\times F$ learnable parameters per level. The resolutions of the hash grids are arranged in a geometric progression: $N_{min}$, $b\times N_{min}$, $\ldots$, $b^{(L-1)}\times N_{min}$, where $N_{min}$ and $b$ denote the initial term and common ratio of the progression, respectively. Coarser grid nodes cover larger coordinate regions, thereby helping to preserve non-local continuity of the signal, while finer grid nodes correspond to local coordinates, enabling the capture of richer high-frequency details. For any given input coordinate $v$, the encoding function retrieves the feature vectors of the corresponding voxel vertices at each resolution level and performs linear interpolation based on the coordinate’s relative position within the voxel. Finally, the interpolated $F$-dimensional feature vectors from all $L$ levels are concatenated to form the final coordinate encoding vector $\phi(v)$ with an output dimension of $L\times F$.
The INR-based neural network is used as the regularization term within each basic unit of the unrolled iterative network. The process is represented as (10)
$z^{t}=f_{\theta}(\phi(v)),$ (10)
where $z^{t}$ represents the intermediate variable generated by the $t$th iteration.
For the data consistency (DC) term within the basic unit, that is, the solution of (7), the conjugate gradient (CG) optimization algorithm is employed. This method inherently requires multiple iterative computations. The DC term comprising multiple CG steps is capable of accurately satisfying the data consistency constraint in each iteration. It is noteworthy that since the CG algorithm itself contains no trainable parameters, backpropagation can be performed without storing intermediate results from each CG step. This characteristic allows a large number of CG steps to be executed in each training iteration with almost no increase in memory usage.
3.3 Loss Function
The loss function is composed of a data fidelity term and a total variation-based regularization term, expressed as:
$L_{total}=L_{DC}+\lambda_{s}L_{TV},$ (11)
where $\lambda_{s}$ is the penalty parameter that balances the two loss terms.
Data consistency with the original undersampled k-space measurements $y$ is ensured by minimizing the data fidelity term $L_{DC}$, which is defined using a normalized $l_{1}-l_{2}$ loss as:
$L_{DC}=\frac{\|y-\widehat{y}\|_{2}}{\|y\|_{2}}+\frac{\|y-\widehat{y}\|_{1}}{\|y\|_{1}},$ (12)
where $\widehat{y}$ is obtained by transforming the final reconstructed output image $\widehat{x}$ of the network into k-space using the forward model $E$, that is, $\widehat{y}=E\widehat{x}$.
Minimizing the total variation (TV) loss $L_{TV}$ can enhance local spatial consistency, eliminate image noise, and preserve edges in the reconstructed image. The TV loss $L_{TV}$ operates on the final reconstructed output image $\widehat{x}$ and is represented as:
$L_{TV}=\|G\widehat{x}\|_{1},$ (13)
where $G$ represents the gradient operator.
4 Experiments
4.1 Datasets
In this study, three datasets were used to evaluate the performance of the proposed method. The first two datasets were utilized for retrospective experimental testing, while the last dataset was employed for prospective experimental testing.
The retrospective experiments utilized publicly available fully-sampled multi-coil knee and brain MRI data from the fastMRI dataset [107]. The fastMRI dataset contains MRI data with various contrast weightings. The knee data consisted of T1-weighted images acquired from 15 receiver coils. All knee images were center-cropped to a size of $368\times 368$. To train the supervised methods, 739 slices from 29 subjects were selected as the training dataset. The brain data consisted of T2-weighted images acquired from 20 receiver coils. All brain images were center-cropped to a size of $320\times 320$. Similarly, for training the supervised methods, 150 slices from 15 subjects were selected as the training dataset.
The prospectively undersampled data were acquired using a 3T scanner (TIM TRIO, Siemens, Erlangen, Germany) equipped with a 12-channel head coil, employing a fast spin-echo sequence with a spin-locking frequency of 500 Hz. The acquired images had a matrix size of $384\times 384$. The study was approved by the local institutional review board, and informed consent was obtained from all volunteers. Similarly, for training the supervised methods, an additional 112 fully-sampled slices from 7 subjects were used as the training dataset.
4.2 Performance Evaluation
The proposed method was compared with several state-of-the-art MRI reconstruction methods, including MoDL [58], ZS-SSL [105], IMJENSE [66], ConvDecoder [64], and L1-ESPIRiT [101]. MoDL is a supervised approach employing an unrolled iterative framework with a convolutional neural network (CNN)-based regularization term [58]. ZS-SSL is a scan-specific self-supervised method based on an unrolled iterative framework, which similarly utilizes a CNN-based regularizer [105]. IMJENSE introduces a scan-specific self-supervised method leveraging an INR-based network to learn a continuous functional representation of MRI images directly from the undersampled $k$-space [66]. ConvDecoder addresses fast MRI reconstruction using a deep image prior (DIP)-based scan-specific approach [64]. L1-ESPIRiT is a traditional parallel MRI reconstruction method employing total variation as the regularization term, implemented using the BART toolkit in our experiments [101, 62].
To comprehensively evaluate the reconstruction performance of these methods, this study adopts a combined quantitative and qualitative evaluation approach. For quantitative assessment, peak signal-to-noise ratio (PSNR) and structural similarity index (SSIM) were selected as evaluation metrics. PSNR focuses on pixel-level error analysis, while SSIM better aligns with the perceptual characteristics of the human visual system. Generally, higher values of PSNR and SSIM indicate better reconstruction quality. The maximum value of SSIM is 1, and values closer to 1 reflect higher consistency between the reconstruction results and the ground truth images. For qualitative assessment, a visual comparison of the reconstruction results from different methods was conducted. To more clearly demonstrate the reconstruction effects, alongside the full reconstructed images, zoomed-in views of key regions and their corresponding error maps are provided to highlight differences in detail reconstruction among the methods.
4.3 Implementation Details
In the experiments, the number of basic units in the unrolled iterative framework of the proposed UnrollINR was set to 1. The multi-resolution hash encoding and MLP within the INR-based regularization term were implemented using the tiny CUDA neural networks [86] library. The data consistency term was solved using the CG method, with the number of internal iterations for this module set to 20. For all datasets used in the experiments, coil sensitivity maps were estimated from the central region of the fully sampled $k$-space using ESPIRiT with default parameters [101].
Regarding the selection of hyperparameters, the proposed UnrollINR involves two hyperparameters: the regularization parameter $\lambda$ in the unrolled framework and the penalty parameter $\lambda_{s}$ in the loss function. These two hyperparameters were set as learnable parameters. They were initialized with specific values at the beginning of the experiment and were automatically updated during the training process. The selection of these initial values will be analyzed in subsequent experiments. For the retrospective experiments, the initial values were set as $\lambda=0.01$ and $\lambda_{s}=0.5$. For the prospective experiments, the initial values were set as $\lambda=0.05$ and $\lambda_{s}=2.0$.
The proposed method was implemented in PyTorch 1.10 using Python 3.9, and all experiments were conducted on a workstation equipped with an NVIDIA A100 GPU (80 GB).
4.4 Ablation Study
We conducted comprehensive ablation studies to evaluate the impact of key components in the proposed UnrollINR method on the reconstruction performance. Various model variants were designed by modifying individual components, with the full framework serving as the baseline. First, the effectiveness of the basic unit was validated. The basic unit integrates a regularization term and a data consistency term; ablation experiments were performed by removing each term individually to assess their contributions to the reconstruction performance. Second, to examine the influence of the coordinate encoding method within the regularization term, the default Instant-NGP [87] encoding was replaced with DINER [111]. Both coordinate encoding methods are learnable and have been widely adopted in various tasks. Finally, the contribution of the loss function was investigated. The loss function consists of a data fidelity term and a regularization term. The effect of the TV regularization term was evaluated by removing it from the loss function.
5 Results
5.1 Retrospective Reconstruction
For the retrospective experiments, random undersampling masks were used to downsample the data. The datasets used in the retrospective experiments, along with the corresponding acceleration rates R, undersampling rates u_rate, and ACS sizes, are summarized in Table 1.
Table 1: Datasets and the Corresponding acceleration rates R, undersampling rates u_rate, and ACS sizes
Datasets R u_rate ACS size
FastMRI
knee dataset
6 16.58% 18
8 12.50% 16
10 9.78% 12
FastMRI
brain dataset
6 16.56% 18
8 12.50% 16
10 10.00% 12
Figure. 2 displays the retrospective reconstruction results of all comparative methods on a randomly selected knee slice under different acceleration rates R. Quantitative evaluation metrics PSNR and SSIM for reconstruction performance are annotated at the bottom of each reconstructed image. Figure. 2 also shows local magnified views of the reconstructed images and their corresponding absolute error maps. On the far left of the figure, the undersampling masks used for the corresponding acceleration rates are displayed. Based on the quantitative metrics, the proposed method, UnrollINR, achieves the best reconstruction performance across all three acceleration rates, significantly outperforming other comparative methods. In terms of visual results, the unsupervised methods ConvDecoder and L1-ESPIRiT exhibit noticeable artifacts and overly smooth outcomes. The images reconstructed using IMJENSE show apparent blurring and loss of fine details. The unrolled ZS-SSL method effectively removes undersampling artifacts but requires further improvement in detail recovery. The supervised algorithm MoDL also exhibits noticeable residual artifacts. In contrast, the proposed method UnrollINR effectively eliminates undersampling artifacts and achieves the recovery of fine details.
Figure 2: Comparative results of different methods on the fastMRI knee dataset under various acceleration rates R. The quantitative metrics PSNR and SSIM are indicated at the bottom of each reconstructed image. Local magnified views of the reconstructed images and absolute error maps are provided. The undersampling masks used for the corresponding acceleration rates R are displayed on the far left of the figure.
Figure. 3 presents the retrospective reconstruction results of all comparative methods on a randomly selected brain slice under different acceleration rates R. Similarly, Figure. 3 provides locally magnified views of the reconstructed images and error maps, along with the undersampling masks used for the corresponding acceleration rates. According to the quantitative evaluation metrics of the reconstructed images, the proposed UnrollINR method again achieves the best reconstruction performance. Visually, the proposed UnrollINR also demonstrates effective artifact removal and fine detail recovery. The supervised method MoDL shows noticeable residual artifacts at high acceleration rates. The unrolled ZS-SSL method loses fine details in its reconstructed images. Other comparative unsupervised methods exhibit significant noise and residual artifacts.
Figure 3: Comparative results of different methods on the fastMRI brain dataset under various acceleration rates R. The quantitative metrics PSNR and SSIM are indicated at the bottom of each reconstructed image. Local magnified views of the reconstructed images and absolute error maps are provided. The undersampling masks used for the corresponding acceleration rates R are displayed on the far left of the figure.
To further validate the performance of the proposed method, retrospective experiments were conducted on 30 randomly selected slices from the fastMRI knee dataset under acceleration rates R=6, 8, and 10. Table 2 summarizes the results of the retrospective experiments, showing the mean and standard deviation of the quantitative evaluation metrics PSNR and SSIM for the 30 slices. The quantitative metrics of all methods decreased as the acceleration rate increased. The proposed method, UnrollINR, demonstrated the best performance among all comparative methods, achieving the highest PSNR and SSIM values across all three acceleration rates. These results are consistent with the aforementioned visual results, further validating the superiority of the proposed UnrollINR method.
Table 2: Summary of reconstruction results for all comparative methods on the fastMRI knee dataset under different acceleration rates R. The values represent the mean and standard deviation of the quantitative evaluation metrics (PSNR and SSIM) across 30 slices, with the optimal results highlighted in bold.
R Metrics Ourrecon MoDL ZS-SSL IMJENSE ConvDecoder L1-ESPIRiT
[…]
5.2 Prospective Reconstruction
Figure. 4 presents the prospective reconstruction results of all comparative methods with an acceleration rate R of 8 and an undersampling rate of 12.5%. To facilitate comparison, locally magnified views of each reconstructed image are provided. The undersampling mask used in the prospective experiment, which has an ACS size of 26, is displayed on the far left of the figure. As shown in the results, the reconstructed images of L1-ESPIRiT and ConvDecoder appear overly smooth; aliasing artifacts persist in the reconstructions of ZS-SSL and IMJENSE; while the MoDL method exhibits image blurring and noise. In contrast, UnrollINR demonstrates superior performance in artifact suppression and detail preservation, highlighting its precise reconstruction capability.
Figure 4: Comparative results of prospective reconstruction for all methods at an acceleration rate R of 8. Locally magnified views of each reconstructed image are provided. The undersampling mask used in the prospective experiment, with an ACS size of 26, is displayed on the leftmost side of the figure.
5.3 Results of Ablation Study
Table 3 summarizes the quantitative evaluation metrics of reconstruction quality under different ablation study settings at acceleration rates R=6 and 10. The ablation experiments were conducted on 10 randomly selected slices from the fastMRI knee dataset, with the mean and standard deviation of the quantitative metrics PSNR and SSIM for these 10 slices statistically reported.
As shown in the results of Table 3, the complete framework of the proposed UnrollINR consistently achieves the highest reconstruction accuracy. The removal of either the regularization term or the data consistency term leads to a significant decline in reconstruction performance, demonstrating the necessity of these key components in the basic unit. The impact of the regularization term is particularly notable, underscoring the importance of its design. To evaluate the influence of the coordinate encoding function, the selected coordinate encoding method was replaced. The multi-resolution hash encoding Instant-NGP adopted in the baseline outperforms the alternative DINER by better leveraging local features, and the experimental results confirm the effectiveness of the coordinate encoding approach in improving reconstruction accuracy. In experiments concerning the loss function, the removal of the TV regularization term also resulted in varying degrees of performance degradation. Furthermore, the results indicate that at higher acceleration rates, these key modules play an even more critical role, affirming their necessity for accurate reconstruction.
Table 3: summary of ablation experiment results on the fastMRI knee dataset under different acceleration rates R. The values represent the mean and standard deviation of the quantitative evaluation metrics (PSNR and SSIM) for 10 slices, with the optimal results highlighted in bold.
R Ablation Variant PSNR SSIM
[…]
6 Discussion
This study proposes a self-supervised deep unrolled network, UnrollINR, to achieve robust and fast parallel MRI reconstruction. The network utilizes INR-based network as a regularization term in the deep unrolled framework, effectively suppressing undersampling aliasing artifacts and restoring image details. Comparative experiments on both retrospective and prospective datasets demonstrate that the proposed UnrollINR significantly outperforms other comparative methods in reconstruction performance, exhibiting considerable potential for further accelerating MRI acquisition.
6.1 Optimization and Adaptive Tuning of Hyperparameters
In terms of hyperparameter tuning, this study involves two key hyperparameters: the regularization parameter $\lambda$ in the unrolled framework, and the penalty parameter $\lambda_{s}$ in the loss function. In the proposed UnrollINR method, both $\lambda$ and $\lambda_{s}$ are designed as learnable parameters that can be dynamically updated during training. To determine their initial values, this study employs the Ray Tune automated hyperparameter optimization framework integrated with Optuna, using the PSNR between the reconstructed images and the ground truth as the optimization metric. The hyperparameter search ranges are set as follows: $\lambda$ from 0.01 to 0.10 with a step size of 0.01, and $\lambda_{s}$ from 0.1 to 1.0 with a step size of 0.1.
To evaluate the sensitivity of the proposed UnrollINR method to hyperparameters, experiments were conducted on the fastMRI knee dataset with an acceleration rate R of 10. In each test, while adjusting one hyperparameter, the value of the other hyperparameter remained fixed. The experimental results are documented in Figure. 5. By selecting appropriate initial values for the hyperparameters and incorporating an adaptive tuning strategy, a balance between data consistency and regularization can be achieved, ensuring robust reconstruction performance.
Figure 5: The impact of two key hyperparameters in UnrollINR on PSNR values using the fastMRI knee dataset with an acceleration rate R of 10.
6.2 Generalizability of UnrollINR to Different Sampling Pattern
This study further investigates the adaptability of the proposed UnrollINR to different undersampling patterns. To comprehensively evaluate the model’s generalization capability, three representative undersampling patterns in MRI were employed: uniform, radial, and spiral. All sampling patterns were set at an acceleration rate of approximately 10. The experimental results, shown in Figure. 6, demonstrate that the proposed UnrollINR maintains consistent and superior reconstruction performance across various sampling patterns, effectively eliminating the distinct aliasing artifacts introduced by different undersampling patterns. These results indicate that the method is insensitive to sampling patterns, exhibiting strong robustness and generalization capability.
Figure 6: summary of reconstruction results on the fastMRI knee dataset under different undersampling patterns. Quantitative metrics (PSNR and SSIM) are indicated at the bottom of each reconstructed image. Absolute error maps between the reconstructed images and the ground truth are also provided. The corresponding undersampling patterns are displayed at the bottom of the figure, with the undersampling rate u_rate and acceleration rate R annotated for each pattern.
6.3 Choice for the Number of CG Iterations and Basic Units
The number of CG iterations refers to the iteration count within the DC term of the basic unit. To investigate the impact of the CG iteration number on reconstruction performance, five iteration counts ranging from 10 to 30 with an interval of 5 were selected. Experiments were conducted on the fastMRI knee dataset with an acceleration rate R of 10, and the results are presented in Figure. 7. The reconstruction performance improves as the number of CG iterations increases. However, the performance gain diminishes with further increments in the iteration count. Notably, the training time increases substantially with higher iteration numbers. Balancing computational cost and reconstruction quality, the optimal CG iteration number was ultimately set to 20.
Figure 7: Impact of the number of CG iterations on quantitative metrics, PSNR and SSIM, using the fastMRI knee dataset at an acceleration rate R = 10.
Similarly, the selection of the number of basic units was determined by balancing training cost and reconstruction performance. During experiments, increasing the number of basic units did not yield significant improvements in reconstruction performance but led to a substantial increase in training cost. Consequently, the number of basic units was ultimately set to 1.
6.4 Effects of the Regularization Term
Physics-guided deep unrolled iterative methods decompose the inverse problem of fast MRI reconstruction into two subproblems: regularization and data consistency. These subproblems are then alternately solved within a fixed number of iterations using optimization algorithms. Current research on various unrolled iterative algorithms primarily differs in the neural network architectures employed for the regularization term and the solver used for the data consistency term. The data consistency subproblem is typically addressed using linear methods such as gradient descent or conjugate gradient, while the regularization subproblem is implicitly solved with different types of neural networks.
Currently, CNN-based networks are commonly employed as regularization terms in unrolled networks, such as ZS-SSL used as a comparative method in this study. Meanwhile, methods based on INR have achieved remarkable progress in MRI reconstruction due to their ability to effectively capture high-frequency signals and their inherent learning-induced consistency bias, which serves as an effective implicit regularization mechanism for solving inverse problems. This study integrates the advantages of unrolled iterative optimization algorithms with INR-based regularization terms, significantly improving MRI reconstruction performance. The results of the aforementioned comparative experiments further validate the superiority of using INR-based networks as regularization terms.
6.5 Limitation and Future Work
Although the proposed UnrollINR has achieved promising results in fast MRI reconstruction, several aspects remain open for future improvement:
  1. 1.
    Compared to other INR-based methods such as IMJENSE, the training process of our approach remains relatively time-consuming. Future work may explore alternative INR-based iterative frameworks to accelerate convergence.
  2. 2.
    The INR-based regularization network in the current framework primarily relies on the mapping between spatial coordinates and image intensity. Subsequent efforts could focus on developing more effective image-domain encoders to incorporate richer prior knowledge, thereby improving both training efficiency and model performance.
  3. 3.
    Further investigation into other acceleration strategies is also warranted. For instance, applying meta-learning to initialize MLP weights for specific types of signals, or leveraging transfer learning to adapt pre-trained weights to new reconstruction tasks. These techniques may facilitate faster convergence and better performance when handling new data.
7 Conclusion
In this study, a novel zero-shot self-supervised reconstruction method named UnrollINR is proposed, which enables high-quality fast MRI reconstruction without the need for fully sampled data. The proposed method is built upon a physics-guided unrolled iterative reconstruction framework and introduces an INR-based network as an effective explicit regularization prior. By effectively integrating unrolled iterative algorithms with INR-based networks, the method significantly enhances MRI reconstruction performance while maintaining model interpretability. Experimental results demonstrate that the proposed method achieves superior reconstruction quality compared to other methods, while maintaining reliable performance even under highly accelerated rates.
{refcontext}
[sorting = none]